# Identification and estimation of vaccine effectiveness in the test-negative design under equi-confounding

### v2

@@ 1 Introduction
Post-market evaluation of vaccine effectiveness is critical for understanding how vaccines perform in real-world settings, particularly when conditions deviate from those that predominated in pre-market randomized trials [patel_postlicensure_2020]. For instance, as population immunity changes and pathogens undergo antigenic mutation or selection, pre-market effectiveness estimates may no longer be valid [hitchings_effectiveness_2021, israel_elapsed_2021]. Investigators may also want to evaluate vaccine performance in certain subpopulations that were either ineligible or underrepresented [olson_effectiveness_2022], assess outcomes and safety signals that may have been underpowered [thompson_effectiveness_2021], or compare vaccine formulations head-to-head or against other therapeutics not considered in the original trials [skowronski_two-dose_2022]. Often, these evaluations must rely on observational data [chua_use_2020-1, dean_covid-19_2021]. Under these circumstances, the test-negative design (TND) is a cost-effective and efficient study design for estimating vaccine effectiveness [sullivan_potential_2014, jackson_test-negative_2013].
In the literature, the term “test-negative design” has sometimes been applied loosely to any study where individuals who test positive for a disease of interest are compared with those who test negative. However, not all such studies are methodologically equivalent. Here, we focus on the canonical design with the following criteria: patients with acute respiratory illness are prospectively recruited using a unified symptom screen through hospitals or other care facilities, consented, and a laboratory test for the pathogen of interest is performed. Detailed information about their demographic, medical, and vaccination history is collected via a questionnaire at enrollment or through linked medical and vaccination records. Those who test positive form the “cases” and those who test negative are the “controls”. Vaccine effectiveness is then estimated via the case status odds ratio among those screened and tested.
This version of the TND has been promoted on the grounds that it is robust to unmeasured confounding by healthcare-seeking behavior — a common source of bias in observational evaluations of vaccines and in particular seasonal influenza vaccines [jackson_test-negative_2013]. Because participants seek a test conditional on the same symptom set but without knowledge of the causative agent, it was claimed the design effectively restricts to care-seekers, thereby precluding it as potential confounder [jackson_test-negative_2013]. It was also suggested that, by limiting to individuals who actually received a test, the TND may be less subject to measurement error [jackson_test-negative_2013]. Subsequent work pointed out that the TND (i) is subject to residual confounding when healthcare-seeking is nonbinary or when there are other factors affecting vaccine uptake and risk of infection [sullivan_theoretical_2016, lewnard_theoretical_2021, lipsitch_observational_2016], (ii) is subject to possible selection bias due to conditioning on the post exposure outcome of receiving a test [sullivan_theoretical_2016, lipsitch_observational_2016], and (iii) is biased when vaccination has a direct effect on testing behavior [foppa_case_2013].
@@
In this article, we return to the original motivation and show that the TND may be justified under a less restrictive alternative assumption that unmeasured confounding is equivalent for test-positive cases and test-negative controls on the odds ratio scale. This assumption builds on the insight that infection with another respiratory illness, if unaffected by focal vaccination, is effectively a negative outcome control [lipsitch_negative_2010, shi_selective_2020]. Therefore any observed differences in the incidence of test negative illnesses between vaccinated and unvaccinated individuals likely reflect residual confounding. Under the assumption of confounding equivalence, we can then use the difference in test-negative illness to de-bias our estimate of vaccine effectiveness, similar to the bias correction approach of the parallel trends assumption in the difference-in-differences literature [sofer_negative_2016, park_universal_2023, tchetgen_universal_2023]. As we show, this approach remains valid under the outcome-dependent sampling of the TND, where only those who are tested are sampled, if we further assume that (potentially unobserved) determinants of care-seeking act equivalently for test-positive and test-negative illnesses, which, as suggested above, could be justified given the causative agent is unknown to individuals prior to receiving a test. When our assumptions hold, we show that the traditional odds ratio estimator from the TND identifies the conditional causal risk ratio among the vaccinated in the population. We also derive new estimators of the marginal risk ratio among the vaccinated, including those based on outcome modeling and inverse probability weighting. We discuss the implications of our assumptions for the design of TNDs and provide proofs of our results as well as simulations to illustrate the consequences when our assumptions are violated. We also propose a sensitivity analysis to evaluate the impact of deviations from equi-confounding on our effect estimates.
@@ 2 Setup and observed data
Let $X$ represent a vector of baseline covariates, $V$ an indicator of vaccination status (1: vaccinated, 0: unvaccinated), and $Y$ a tri-level outcome at the end of follow up where
$Y:=\begin{cases}Y=2&\text{test positive case},\\ Y=1&\text{test negative case},\\ Y=0&\text{when untested}.\end{cases}$
For simplicity, we consider the canonical TND setting in which a vaccination campaign occurs prior to a seasonal outbreak, allowing vaccination to be a point treatment.
The outcome $Y$ is only meaningful with additional context about the testing regime in the population of interest. We define two intermediate outcomes that define the assumed testing regime in a TND and thus jointly determine $Y$. First, let $I$ denote a categorical indicator of symptomatic infection (2: symptoms caused by pathogen of interest, 1: symptoms caused by something else, 0: no symptomatic illness), where “symptomatic” means the presence of the pre-specified symptom set used to screen for inclusion in the TND study. Note, we assume that the chance of co-infection or repeated infections of different origin are negligible over the follow up interval such that $I=1$ and $I=2$ are mutually exclusive events. Second, let $T$ denote an indicator of seeking a test for the pathogen of interest at an eligible health facility (1: tested, 0: not tested). Combining the two, we assume (i) that participants are screened for symptoms prior to testing such that $T_{i}=1\implies I_{i}\neq 0$ for every individual $i$, (ii) that everyone screened for symptoms is tested, and (iii) that the test performed is perfect, although, as we show, the second and third conditions can be relaxed. Under this testing regime, we have that $Y=I\times T$ and we consider $Y\neq 0$ to define medically-attended illness, i.e., symptomatic infection severe enough to warrant medical attention, with $Y=2$ denoting medically-attended illness caused by the focal pathogen targeted by the vaccine and $Y=1$ denoting medically-attended illness due to something else. Finally, throughout we denote by $Y^{v}$ the potential outcome under an intervention that sets $V=v$ for $v=0,1$.
Suppose we have data from a TND study in which $n$ participants are tested under the regime specified above. Denoting selection into the TND study by $S$ and using the definitions above, the data are assumed to be independent realizations
$O_{TND}=\{(X_{i},V_{i},S_{i}=1,Y_{i}):i=1,\ldots,n\},$
from the law of $(X,V,Y)\mid S=1$. Under the assumed regime, $S=\mathbbm{1}(Y\neq 0)$ as data are only recorded among those screened and tested. However, our setup could be extended to other testing regimes such as first requiring hospitalization prior to testing or incorporating refusals among those offered a test.
At points, it will be useful to consider, as an alternative to TND sampling, a design in which $N$ individuals are recruited at baseline and followed forward, as in a cohort study, through the same screening and testing regime such that the data are instead independent realizations of
$O_{cohort}=\{(X_{i},V_{i},Y_{i}):i=1,\ldots,N\}$
from the law of $(X,V,Y)$. Note that, under this sampling design, we could still recover the data observed in the TND, assuming equivalent testing and screening regime, by subsetting to individuals who receive a test but the reverse, recovering data observed under cohort sampling from a TND, will generally not be possible.
@@ 3 Causal estimands for the test-negative design
Traditionally, in a TND study, the target parameter is the causal risk ratio,
$\Psi_{RR}\equiv\dfrac{\Pr[Y^{1}=2]}{\Pr[Y^{0}=2]},$
which contrasts the probability of medically-attended illness due to the pathogen of interest if everyone were vaccinated compared to if they were unvaccinated in the population underlying the TND, with causal vaccine effectiveness (VE) defined as $1-\Psi_{RR}$. In previous literature, the causal odds ratio is sometimes presented as the target parameter instead of the risk ratio. However, this is often combined with an assumption —either explicit or implicit— that symptomatic infection is rare in the population underlying the TND over the follow up period, in which case the odds ratio approximates the risk ratio.
@@
may nevertheless be point identified under reasonable conditions on the risk ratio scale. The parameter $\Psi$ is analogous to the average treatment effect on the treated (ATT) in the causal inference literature, though expressed as a ratio rather than a difference. In light of this observation, we redefine causal VE can therefore be defined as $1-\Psi$, which can be interpreted as the proportion of medically-attended illness prevented by vaccination among the vaccinated subjects. When vaccination provides constant protection for everyone, $\Psi$ will coincide with the causal risk ratio in the full population, i.e. $\Psi_{RR}=\Psi$. However, when protection is heterogeneous, the causal risk ratio in the full population becomes a weighted average of the causal risk ratio among the vaccinated and unvaccinated.
A related estimand of interest under effect heterogeneity is the conditional causal risk ratio among the vaccinated, i.e.
@@
While $\Psi$ is our preferred estimand, we introduce $\Psi(X)$ for two reasons (1) it is the estimand targeted by the traditional logistic regression estimator in a TND study when effect homogeneity is assumed or when heterogeneity is modeled via the inclusion of product terms and (2) it allows us to introduce our main identifiability arguments in simpler form.
@@ 4 Illustrative example
To build intuition, we start with a brief account of the identification of the conditional risk ratio among the vaccinated, $\Psi(X)$, before introducing our formal identification results for $\Psi$ in section 5.
@@ 4.1 Decomposition of the conditional causal risk ratio among the vaccinated
Consider the estimand $\Psi(X)$. Multiplying by unity in the form of $\frac{\Pr[Y^{0}=2|V=0,X]}{\Pr[Y^{0}=2|V=0,X]}$ and applying consistency of potential outcomes (Assumption A1 below) leads to the following decomposition:
@@
This equation shows $\Psi(X)$ can be written as the product of two terms: the (crude) observed risk ratio and a de-biasing term. The de-biasing term represents the ratio of the potential outcomes if vaccination were withheld between those who were, in fact, vaccinated and those who were not. It quantifies the degree of unmeasured confounding between the vaccinated and unvaccinated on the ratio scale, i.e., the net influence of all unmeasured causes of vaccination and the outcome. When there is no unmeasured confounding, the de-biasing term equals one and the observed risk ratio equals the causal risk ratio among the vaccinated. Conversely, in the presence of unmeasured confounding, the de-biasing term is the quantity necessary to remove bias from the observed risk ratio. However, without additional assumptions, this term (and in particular the denominator) cannot be identified from the observed data.
@@ 4.2 The test-negative illness as a proxy for unmeasured confounding
One approach to identifying the de-biasing term, $\frac{\Pr[Y^{0}=2|V=0,X]}{\Pr[Y^{0}=2|V=1,X]}$, is to find a proxy that is observed and could therefore be used for bias correction. A natural proxy is another related infection (for instance one producing the same symptoms) that shares the same hidden factors confounding the relationship between vaccination and focal infection and testing, but is itself unaffected by vaccination. In a TND this is the role of the test-negative illness. By design, those testing negative in the TND demonstrate a similar propensity for infection and for seeking care when symptomatic. Assuming the vaccine has no effect on the test negative illness (Assumption A2 below), any observed difference in the incidence of test-negative illness between the vaccinated and unvaccinated, after conditioning on measured covariates $X$, must reflect residual unmeasured confounding. When the ratio of test-negative illness among the vaccinated and unvaccinated is a perfect proxy for the de-biasing term, i.e., when
$\dfrac{\Pr[Y^{0}=2|V=1,X]}{\Pr[Y^{0}=2|V=0,X]}=\frac{\Pr[Y=1|V=1,X]}{\Pr[Y=1|V=0,X]}.$ (2)
we call this odds ratio equi-confounding (Assumption A3 below) as it implies that any unmeasured factors $U$ exert identical effects on both test-positive and test-negative illnesses on the odds ratio scale (Note in this case, because the illnesses are also assumed to be mutually exclusive, the odds ratio scale is also the risk ratio scale).
@@ 4.3 Identification under cohort and outcome-dependent sampling
Substituting the ratio of test-negative illness in the vaccinated and unvaccinated for the de-biasing term in our original decomposition of $\Psi(X)$, we derive the following expression:
@@ 5 Identification
We now return to the marginal causal risk ratio among the vaccinated, $\Psi$, and provide formal identifiability conditions and results which follow the same logic as our account of the identification of $\Psi(X)$ above.
@@ 5.1 Identifiability conditions
Assumption A1 is a well-known identifiability condition in the causal inference literature discussed in more detail elsewhere [hernan_causal_2020]. Assumption A2 has two implications. First, it implies that the vaccine does not offer any cross-protection against other types of infection which may cause the same symptoms, including any short-term, nonspecific protection via activation of the immune system. Second, it requires that an individual’s decision to seek care and get tested is unaffected by their vaccination status, though it may depend on other measured and unmeasured factors. In theory, Assumption A2 could be evaluated in a randomized trial by comparing the incidence of testing negative (or testing positive for specific non-target pathogens) across the vaccine and comparison arms. However, in practice, trials may not be adequately powered to rule out small effects. In placebo-controlled trials, differences in care-seeking are minimized through the blinding of the participant to their vaccination status. Both aspects of Assumption A2 have been invoked in previous literature on analysis of TND studies [jackson_test-negative_2013, feng_assessment_2017, schnitzer_estimands_2022]. In Appendix F, we show that Assumption A2 can be further relaxed if the effect of vaccination on testing is equivalent for test-positive and test-negative illnesses.
@@
Assumptions A3a and A3b are technically stronger than A3 alone as A3 allows, in principle, for situations in which the terms in Assumptions A3a and A3b are both unequal but they cancel when multiplied such that Assumption A3 still holds.
Finally, Assumption A4 is similar to, although less restrictive than, the well-known positivity condition in causal inference [hernan_causal_2020]. Specifically, Assumption A4 requires overlap of vaccination propensity between the observed vaccinated and unvaccinated among the test positive cases as well as overlap of vaccination propensity across test-positive cases and test-negative controls within observed vaccination groups. Taken together, the conditions $\mathcal{S}_{2}(1)\subseteq\mathcal{S}_{2}(0)$ and $\mathcal{S}_{2}(v)\subseteq\mathcal{S}_{1}(v)$ ensure that the odds ratios $OR_{2}(X)$ and $OR_{1}(X)$ are well-defined (i.e. nonzero in numerator and denominator).

### v3

@@ 1 Introduction
Post-market evaluation of vaccine effectiveness is critical for understanding how vaccines perform in real-world settings, especially as population immunity changes or pathogens undergo antigenic mutation or selection [patel_postlicensure_2020, hitchings_effectiveness_2021, israel_elapsed_2021]. These studies can also assess effectiveness in subpopulations that were ineligible or underrepresented [olson_effectiveness_2022], evaluate outcomes and safety signals that may have been underpowered [thompson_effectiveness_2021], or compare vaccine formulations not considered in the original trials [skowronski_two-dose_2022]. Often, these evaluations must rely on observational data [chua_use_2020-1, dean_covid-19_2021].
The test-negative design (TND) is a common and cost-efficient study design for estimating vaccine effectiveness [sullivan_potential_2014, jackson_test-negative_2013]. In its canonical form, patients with acute respiratory illness are prospectively screened and recruited based on a unified symptom set through hospitals or other care facilities and tested for the pathogen of interest. Vaccination history and other detailed covariate information are collected at enrollment or through linked medical or vaccination records. Effectiveness is then estimated by comparing the odds of vaccination among test-positive (cases) and test-negative (controls) individuals.
The TND has been previously promoted as a way to eliminate bias due to unmeasured healthcare-seeking behavior — a common source of bias in vaccine evaluations [jackson_test-negative_2013]. Because participants seek a test conditional on the same symptom set but without knowledge of the causative agent, the design effectively restricts to “care-seekers”, thereby precluding it as potential confounder [jackson_test-negative_2013]. Subsequent work pointed out that the TND (i) may not be robust to unmeasured healthcare-seeking when it is nonbinary or when there are other factors affecting vaccine uptake and risk of infection [sullivan_theoretical_2016, lewnard_theoretical_2021, lipsitch_observational_2016] and (ii) is subject to possible selection bias due to conditioning on receiving a test [sullivan_theoretical_2016, lipsitch_observational_2016].
@@
In this article, we revisit the original motivation and show that the TND may be justified under a less restrictive assumption that unmeasured confounding is equivalent for test-positive cases and test-negative controls on the odds ratio scale. This assumption builds on the insight that infection with another respiratory illness, if unaffected by focal vaccination, is essentially a negative outcome control [lipsitch_negative_2010, shi_selective_2020]. Therefore, any observed difference in the incidence of test negative illnesses between vaccinated and unvaccinated individuals likely reflects residual confounding. Under the assumption of confounding equivalence, we can use the difference in test-negative illness to de-bias our estimate of vaccine effectiveness, similar to the bias correction via the parallel trends assumption in the difference-in-differences literature [sofer_negative_2016, park_universal_2023, tchetgen_universal_2023]. As we show, this approach is valid under the outcome-dependent sampling of the TND, where only those who are tested are sampled. When our assumptions hold, we show that the traditional odds ratio estimator identifies the conditional causal risk ratio among the vaccinated. We also derive new estimators of the marginal risk ratio among the vaccinated, including those based on outcome modeling and inverse probability weighting. We discuss the implications of our assumptions for the design of TNDs and provide proofs of our results as well as simulations to illustrate the consequences when our assumptions are violated. We also propose a sensitivity analysis to evaluate the impact of deviations from equi-confounding on our effect estimates.
@@ 2 Setup and observed data
Let $X$ represent a vector of baseline covariates, $V$ an indicator of vaccination status (1: vaccinated, 0: unvaccinated), and $Y$ a tri-level outcome at the end of follow up where:
$Y:=\begin{cases}Y=2&\text{test positive case},\\ Y=1&\text{test negative case},\\ Y=0&\text{not tested}.\end{cases}$
For simplicity, we consider the original TND setting in which a vaccination campaign occurs prior to a seasonal outbreak, and thus vaccination is a point treatment.
The outcome $Y$ is only meaningful with additional context about the testing regime in the population of interest. We introduce two additional variables that jointly define this regime. First, a categorical indicator $I$ of symptomatic infection (2: caused by pathogen of interest, 1: other cause, 0: no symptoms), where symptoms are the pre-specified set used to screen for inclusion in the TND study. Second, an indicator of receiving a test at a health facility $T$ (1: tested, 0: not tested).
We further assume:
  1. (i)
    The chance of co-infection or repeated infections of different origin are negligible over the follow up interval such that $I=1$ and $I=2$ are mutually exclusive events;
  2. (ii)
    individuals are screened for symptoms prior to testing such that $T_{i}=1\implies I_{i}\neq 0$ for every individual $i$;
  3. (iii)
    all screened individuals are tested;
  4. (iv)
    the test is perfect
These assumptions can be relaxed (e.g., we discuss imperfect testing in Appendix H). Under this testing regime, we have that $Y=I\times T$ and we consider $Y\neq 0$ to define medically-attended illness, i.e., symptomatic infection severe enough to warrant medical attention. Finally, let $Y^{v}$ denote the potential outcome under $V=v$ for $v=0,1$.
The observed TND data are $n$ independent realizations
$O_{\text{TND}}=\{(X_{i},V_{i},S_{i}=1,Y_{i})\}_{i=1}^{n},$
from the law of $(X,V,Y)\mid S=1$ where $S$ denotes selection into the TND study with $S=\mathbbm{1}(Y\neq 0)$ under the testing regime.
For comparison, we also define data from a hypothetical cohort, i.e. $N$ independent realizations
$O_{\text{cohort}}=\{(X_{i},V_{i},Y_{i})\}_{i=1}^{N},$
from the law of $(X,V,Y)$, under the same screening and testing regime. TND data can be recovered by restricting the cohort to those tested, but not vice versa.
@@ 3 Causal estimands for the test-negative design
In a TND study, the primary estimand is the causal risk ratio,
$RR\equiv\dfrac{\Pr[Y^{1}=2]}{\Pr[Y^{0}=2]},$
comparing the probability of medically-attended illness due to the pathogen of interest if everyone were vaccinated versus unvaccinated in the population underlying the TND, with causal vaccine effectiveness (VE) defined as $1-RR$. Occasionally, the causal odds ratio is presented as the target parameter instead of the risk ratio. However, this is often combined with an assumption —either explicit or implicit— that symptomatic infection is rare, in which case the odds ratio approximates the risk ratio.
@@
may nevertheless be point identified. The parameter $\Psi$ is analogous to the average treatment effect on the treated (ATT) in the causal inference literature, though expressed as a ratio rather than a difference. Similarly, we can define a modified VE as $1-\Psi$, interpreted as the proportion of medically-attended illness prevented by vaccination among the vaccinated subjects. When protection is homogeneous, $\Psi$ coincides with the causal risk ratio, i.e. $RR=\Psi$; otherwise, when protection is heterogeneous, $RR$ is a weighted average of the causal risk ratios among the vaccinated and unvaccinated.
We also consider the conditional causal risk ratio among the vaccinated, i.e.
@@
which is the estimand targeted by traditional logistic regression estimator in a TND study when effect homogeneity is assumed or when heterogeneity is modeled via inclusion of product terms. While $\Psi$ is our preferred estimand, we use $\Psi(X)$ to first present our identification arguments in simplified form.
@@ 4 Illustrative example
To build intuition, we start with a brief account of the identification of the conditional risk ratio among the vaccinated, $\Psi(X)$, before turning to formal identification results for $\Psi$ in section 5.
@@ 4.1 Decomposition of the conditional causal risk ratio among the vaccinated
Consider the estimand $\Psi(X)$. Multiplying by unity in the form of $\frac{\Pr[Y^{0}=2|V=0,X]}{\Pr[Y^{0}=2|V=0,X]}$ and applying consistency of potential outcomes (Assumption A1 below), we can write:
@@
This decomposition expresses $\Psi(X)$ as the product of two terms: the (crude) observed risk ratio and a de-biasing term. The de-biasing term quantifies the degree of unmeasured confounding between the vaccinated and unvaccinated on the ratio scale, i.e., the net influence of all unmeasured causes of vaccination and the outcome. When there is no unmeasured confounding, the de-biasing term equals one and the observed risk ratio equals the causal risk ratio among the vaccinated. Conversely, in the presence of unmeasured confounding, the de-biasing term is the quantity necessary to remove bias from the observed risk ratio. However, without additional assumptions, this term (and in particular the denominator) cannot be identified from the observed data.
@@ 4.2 The test-negative illness as a proxy for unmeasured confounding
To identify the de-biasing term, $\frac{\Pr[Y^{0}=2|V=0,X]}{\Pr[Y^{0}=2|V=1,X]}$, we seek a proxy outcome that could be used in place of test-positive illness for bias correction. A natural choice is another infection that produces the same symptoms and shares many of the same confounders, but is itself unaffected by focal vaccination. In a TND this is the role of the test-negative illness. When the net effect of all unmeasured factors $U$ is equivalent on the ratio scale for test-positive and test-negative illness after conditioning on covariates $X$, i.e. when
$\dfrac{\Pr[Y^{0}=2|V=1,X]}{\Pr[Y^{0}=2|V=0,X]}=\frac{\Pr[Y^{0}=1|V=1,X]}{\Pr[Y^{0}=1|V=0,X]},$ (2)
we call this odds ratio equi-confounding (Assumption A3 below; Note in this case, because the illnesses are also assumed to be mutually exclusive, the odds ratio scale is also the risk ratio scale). Furthermore, if vaccination does not affect the test-negative outcome then
$\frac{\Pr[Y^{0}=1|V=1,X]}{\Pr[Y^{0}=1|V=0,X]}=\frac{\Pr[Y=1|V=1,X]}{\Pr[Y=1|V=0,X]},$ (3)
implying that the de-biasing term is identified by the ratio of test-negative outcome comparing the vaccinated to unvaccinated.
@@ 4.3 Identification under cohort and outcome-dependent sampling
Substituting $\frac{\Pr[Y=1|V=1,X]}{\Pr[Y=1|V=0,X]}$ for the de-biasing term in our original decomposition of $\Psi(X)$, we derive the following expression:
@@ 5 Identification
We now return to the marginal causal risk ratio among the vaccinated, $\Psi$, and formalize our identifiability conditions.
@@ 5.1 Identifiability conditions
Assumption A1 is a well-known identifiability condition discussed in more detail elsewhere [hernan_causal_2020]. Assumption A2 has two implications. First, it implies that the vaccine does not offer any cross-protection against other types of infection which may cause the same symptoms, including any short-term, nonspecific protection via activation of the immune system. Second, it requires that an individual’s decision to seek care and get tested is unaffected by their vaccination status, though it may depend on other measured and unmeasured factors. In theory, Assumption A2 could be evaluated in a randomized trial by comparing the incidence of testing negative (or testing positive for specific non-target pathogens) across arms. However, in practice, trials may not be adequately powered to rule out small effects. Both aspects of Assumption A2 have been invoked in previous literature on analysis of TND studies [jackson_test-negative_2013, feng_assessment_2017, schnitzer_estimands_2022]. In Appendix F, we show that Assumption A2 can be further relaxed if the effect of vaccination on testing is equivalent for test-positive and test-negative illnesses.
@@
Assumptions A3a and A3b are technically stronger than A3 alone as A3 allows, in principle, for situations in which the terms in Assumptions A3a and A3b are both unequal but cancel when multiplied.
Finally, Assumption A4 is similar to, although less restrictive than, the well-known positivity condition in causal inference [hernan_causal_2020]. Specifically, Assumption A4 requires overlap of vaccination propensity between the observed vaccinated and unvaccinated among the test positive cases as well as overlap of vaccination propensity across test-positive cases and test-negative controls within observed vaccination groups.
The causal directed acyclic graphs (DAG) in Figure 1 shows a possible mechanism satisfying the identifiability conditions above.